Express as cm using decimals
9 cm 8 mm
 Express as cm using decimals
9 cm 8 mm

Grade:12

2 Answers

Harshit Singh
2 years ago
Dear Student

We know that there are 10 mm in 1 cm
9 cm 8 mm = [(9 + 8 / 10)] cm
= 9.8 cm

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We know that,
 
10 mm = 1cm
 
1 mm = 1/10 cm
 
8 mm = 8/10 = 0.8 cm
 
9 cm 8 mm = (9+0.8) = 9.8 cm